# When Should Selfish Miners Double-Spend?

### v3

@@ 3.1 Propagation Delays
We note that, from the perspective of the propagation delays, the model described here has been considered in the majority of the selfish mining literature and is still the convention. Essentially, if blocks can be mined during propagation delays, the analysis has to keep track of multiple (potentially infinite) different honest miner views that evolve as time goes by. Hence, in the selfish mining literature, the assumption of no block mining during propagation delays has been adopted for tractability purposes. On the other hand, studies on double-spending analysis, such as [43, 44, 45], consider the propagation delays, because from the perspective of the consistency guarantees, worst-case scenarios are relevant. Hence, these studies usually assume that the adversary controls the propagation delays as long as each propagation delay is lower than a maximum $\Delta$ and ties are broken in its favor. As a result, these analyses usually focus on a single worst-case honest miner view [44, Section III.A, weakest possible view]. However, such a situation is extremely pessimistic as our goal is not to give worst-case analysis of revenue ratios in selfish mining. We also note that, optimal selfish mining analysis has been done previously in such a pessimistic setting [46, Lemma 4].
@@
Further, under Nakamoto consensus, given the block creation rate $\lambda$, if $\lambda\Delta$ is not small, the consistency of the system at threat and selfish mining is a less relevant concern. In such systems, the research focuses on different consensus models such as [47, 48, 49]. On the other hand, as long as $\lambda\Delta$ is small, it can be shown that the probability of more than $2$ blocks being mined within a $\Delta$ interval is negligible [50, 31, 51]. Thus, in the realistic scenario of one block being mined during the propagation delay of a block, during forks, the honest network divides between two branches. As a result, in the long run, the parameter $\gamma$ used in selfish mining literature is still a relevant and reasonable assumption. Further, in [50, Section V] the authors give a formula to estimate $\gamma$ given propagation delays between mining entities and validate the estimation with network simulations.
@@
Essentially, as things become immediately untractable when one considers multiple miners’ views during network delays, the system model with $\alpha$ and $\gamma$ serve as tractable and efficient parameters to model selfish mining attacks in a simplified manner that still roughly translates to a system with propagation delays where $\gamma$ is the main parameter of interest from the perspective of the delays and fork conflicts and should be estimated based on the observed delays. Here, note that, propagation delays also affect the apparent hashrate of a mining entity. In other words, the apparent hashrate might deviate from the fraction of computational power the entity possesses due to the propagation delays and network structure even if everyone follows honest mining strategy. Notice that, propagation delays result in forks and the block published first does not always win the fork race. Hence, one can also think of $\alpha$ as the apparent hash rate instead of the fraction of computational power. For example, studies such as [51] consider different propagation delays between different miners to evaluate fairness, which is roughly defined as the difference between the fraction of computational power and the apparent hashrate a mining entity has. We believe that modeling such a system to estimate $\alpha$ and $\gamma$ values used in this paper are possible however out of scope of this paper. For further details about the propagation delays and their impact on fork races and fairness, we refer the reader to [50, 31, 51].
@@ 3.2 $L$-Stubborn Mining
The selfish mining strategies are extensively analyzed in the literature with notable works such as [4, 6, 5]. In this work, we consider a sub-group of strategies from all possible states and action space, which we call $L$-stubborn [^1] In this paper, $L$-stubborn mining is different than the lead-stubbornness (L-stubbornness) defined in [6]. There, “L” stands for the word “lead.” Here, $L$ refers to an integer which is the length of the adversarial chain with respect to a common ancestor block between the honest and the adversarial chains. mining. To describe the strategy, let us assume that all participants start with the same chain, which we call the offset chain. For the sake of simplicity, we denote the height of the offset chain as zero. Both the adversary and the honest miners mine on top of the offset chain.
@@
    If the adversary is the first to mine a block on top of the offset chain, i.e., at height $1$, it keeps the block private and continues to mine privately on top of its block. We denote the highest block height of the adversarial chain as $A$ and call the adversarial chain as $A$-chain. At any time later, if the honest miners mine a block on at height $h$, the adversary publishes its block at the same height immediately if $h\leq A<L$, which we call matching. In accordance with the network model defined earlier, $\gamma$ proportion of the honest miners accept and mine on top of the adversarial block instead of the honest block. We denote the highest honest block height as $H$ and the chain starting from the offset and ending in the honest block at height $H$ as $H$-chain. The private stubborn mining strategy of the adversary continues on for all heights until $A=L$, i.e., whenever $L>A\geq H$ adversary immediately matches the block at the height of the newly mined honest block.
@@
An attack cycle is a process that starts when an offset chain is defined, and ends when a new offset chain is defined. Note that, offset chain is redefined whenever an adopt or override action (event) happens, which we call an offset event in general.
@@
Note that, for $L=1$ the strategy reduces to honest mining strategy, whereas for $L=2$ the strategy reduces to the celebrated selfish mining attack of [4]. On the other hand, for $L=\infty$ the strategy becomes the equal fork stubborn mining strategy described in [6] and analyzed in closed-form with Catalan numbers in [37]. In this paper, we are interested in giving a method for calculating the revenue ratios for any given $L$ as well as deriving a method to find the best possible $L$ given $\alpha$ and $\gamma$. The resulting strategy does not allow any trail stubbornness discussed in [6], i.e., the adversary accepts the honest chain whenever it falls behind. Hence, we can only claim the optimality of the strategy under action spaces with independent attack cycles that allow no-trail stubbornness. We argue that the parameter $L$ represents the level of stubbornness in selfish mining and our analysis provides the optimal level of stubbornness as well as the maximum stubbornness such that the attack is still more profitable than honest mining.
@@
In [5], the authors use an undiscounted average reward MDP to provide $\epsilon$-optimal policies. The authors of [5] further argue about combining double-spending attack with selfish mining attack, however, they do not provide any concrete method or analysis. Later, in [8], the same authors provide an MDP that accounts for rewards of double-spending attack, however, the MDP ignores rewards from adversarial blocks that do not result in double-spending. Thus, the optimal policies resulting from the MDP therein are not optimal in the sense that they provide maximum mining rewards with free double-spending. Further, the results and strategies provided in [8, 5, 7] are complicated with no easy-to-follow structure. $L$-stubborn mining that we consider in this paper is a policy which results in higher mining rewards than honest mining (as well as optimal ones) and has strong implications on double-spending strategies with a clear and easy-to-follow analysis under $k$-confirmation rule, which we explain next.
3.3 $k$-Confirmation Rule
@@
First, we rigorously define the events related to a double-spending attack.
@@
Consider a transaction $tx$ where the adversary pays a merchant to get a service and a conflicting transaction $tx^{\prime}$ where the adversary simply changes the address of the same funds, which it can use again later. We consider three disjoint events regarding $tx$ and $tx^{\prime}$:
@@
    Service If $tx$ is confirmed and makes it into the offset chain, we call the event Service.
@@
    Move-Funds If $tx$ is never confirmed, we call the event Move-Funds.
@@
    Double-Spending If $tx$ is confirmed but does not make it into the offset chain, we call the event Double-Spending.
Notice the relation between $L$-stubborn mining and $k$-confirmation rule. If $(k+1)$-stubborn mining attack is more profitable than honest mining, adversary can try to double-spend at no-cost [^2] As observed by [5], $tx$ is a real transaction and its confirmation results in a service, therefore, this should not be seen as a loss even from the perspective of the adversary. while implementing the $(k+1)$-stubborn mining strategy. All it needs to do is to include a $tx^{\prime}$ in the first height above the offset chain, whereas honest miners will include $tx$ on the first block above the offset chain. Later on, when $A$-chain is released, if $H$-chain contains $tx$, it will be replaced by $tx^{\prime}$, allowing the adversary to double-spend.
Before the offset event happens, if $A>k$, this implies that at some point $H=k$ and adversary will release $A$-chain later on. Depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Double-Spending happens. On the other hand, before the offset event happens, if $A<H\leq k$, depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Service happens. If an attack cycle ends with Move-Funds event, later, adversary can try double-spending again using the same funds. One could also rescale the probabilities of Double-Spending and Service by excluding Move-Funds event from the perspective of the actual funds associated with $tx$ and $tx^{\prime}$ since eventually the fund is going to result in Service or Double-Spending. We omit the rescaling in our calculations and numerical results, as rescaling is straightforward to do and we would like to present the exact probability of Double-Spending event in each attack cycle.
We also note that the time $tx$ is released in the network is another important issue while considering double-spending attacks. In general, it can be shown that a double-spending attack can be successful with probability $1$ if the adversary is allowed to pick the time $tx$ is released, since it can build an arbitrarily high pre-mining gain [8]. The pre-mining gain refers to the difference between the private chain of the adversary and the public honest chain, in general [8]. The approach we take here is more conservative and assumes that the $tx$ is released at the start of each attack cycle, hence the pre-mining gain is 0.
@@ 3.4 $S$-Stealth Mining
Notice the difference between Nakamoto private attack and $(k+1)$-stubborn mining. The first major difference is $(k+1)$-stubborn mining gives up the attack when $A$-chain falls behind $H$-chain, i.e., no-trailing behind is allowed. Another major difference is, during a cycle, $(k+1)$-stubborn mining matches each honest block with an adversarial one which can expose the attack as two chains with conflicting transactions $tx$ and $tx^{\prime}$ are competing openly in the network, arousing the merchant’s suspicion. The matching also decreases Double-Spending probability while increasing Move-Funds probability, which we show later.
We address some of these issues by modifying $L$-stubborn mining in a way that only allows matching if both double-spending and the risk of losing mining revenue are imminent, i.e., H=A=k=L1H=A=k=L-1. We call this attack $S$-stealth mining, which we investigate in Section 5. Essentially, $S$-stealth mining strategy is the same strategy as $L$-stubborn mining strategy except that the matching action only happens when H=A=S1H=A=S-1. We note that $S$-stealth mining does not allow any trailing behind either, hence, there is still a difference between a private attack and $S$-stealth mining. Note that, similar to $L$ of $L$-stubborn mining strategy, we can view $S$ of $S$-stealth mining strategy as a parameter of stubbornness.

### v4

@@ 2.1 Propagation Delays
Notice, when a block is mined by an honest miner, during the propagation delay, the adversary can rush to release its already mined private blocks but it is inherently assumed that no new blocks are mined during this propagation. Hence, all honest blocks as well as the blocks released by the adversary are treated as public block in this model, i.e., when an honest miner mines a block, all other honest miners also become aware of the block immediately although the adversary can rush and release a block at the same height to influence $\gamma$ fraction of the honest miners who then prefer the adversarial branch. As a result, in this model, honest miners do not fork each other’s blocks. From the perspective of the propagation delays, this model has been used in the majority of the selfish mining literature and is still the convention [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 37, 25, 41, 42, 43, 23, 24, 26].
@@
Essentially, if blocks can be mined during propagation delays, the analysis has to keep track of multiple (potentially infinite) different honest miner views that evolve as time goes by. Hence, in the selfish mining literature, the assumption of no block mining during propagation delays has been adopted for tractability purposes. On the other hand, studies on double-spending analysis, such as [44, 45, 46], consider adversarial blocks mined during the propagation delays, because from the perspective of the consistency guarantees, worst-case scenarios are relevant. Hence, these studies usually assume that the adversary controls the propagation delays as long as each propagation delay is lower than a maximum $\Delta$ and ties are broken in its favor. As a result, these analyses usually focus on a single worst-case honest miner view [45, Section III.A, weakest possible view]. However, such a situation is extremely pessimistic as our goal is not to give worst-case analysis of revenue ratios in selfish mining, rather a realistic evidence that the adversary is incentivized to deviate from honest protocol. We also note that, optimal selfish mining analysis has been done previously in such a pessimistic setting [47, Lemma 4]. The results therein follow an exponentially bounded delay structure, however, they are easily extendable to bounded $\Delta$ delay model of [44], i.e., by replacing $P^{\prime}$ matrix of [47] with $P$ matrix of [46] and redoing the subsequent analysis of [47, Lemma 4].
@@
Further, under Nakamoto consensus, given the block creation rate $\lambda$, if $\lambda\Delta$ is not small, the consistency of the system at threat and selfish mining is a less relevant concern. In such systems, the research focuses on different consensus models such as [48, 49, 50]. On the other hand, as long as $\lambda\Delta$ is small, it can be shown that the probability of more than $2$ blocks being mined within a $\Delta$ interval is negligible [51, 31, 52]. Thus, in the realistic scenario of one block being mined during the propagation delay of a block, during forks, the honest network divides between two branches. As a result, in the long run, the parameter $\gamma$ used in selfish mining literature is still a relevant and reasonable assumption. Further, in [51, Section V] the authors give a formula to estimate $\gamma$ given propagation delays between mining entities and validate the estimation with network simulations.
@@
Essentially, as things become immediately untractable when one considers multiple miners’ views during network delays, the system model with $\alpha$ and $\gamma$ serve as tractable and efficient parameters to model selfish mining attacks in a simplified manner that still roughly translates to a system with propagation delays where $\gamma$ is the main parameter of interest from the perspective of the delays and fork conflicts and should be estimated based on the observed delays. Here, note that, propagation delays also affect the apparent hashrate of a mining entity. In other words, the apparent hashrate might deviate from the fraction of computational power the entity possesses due to the propagation delays and network structure even if everyone follows honest mining strategy. Notice that, propagation delays result in forks and the block released and observed first by other miners does not always win the fork race. Hence, one can also think of $\alpha$ as the apparent hash rate instead of the fraction of computational power. For example, studies such as [52] consider different propagation delays between different miners to evaluate fairness, which is roughly defined as the difference between the fraction of computational power and the apparent hashrate a mining entity has. We believe that modeling such a system to estimate $\alpha$ and $\gamma$ values used in this paper are possible however out of scope of this paper. For further details about the propagation delays and their impact on fork races and fairness, we refer the reader to [51, 31, 52].
3 Mining Strategies and Definitions
@@ 3.1 $L$-Stubborn Mining
The selfish mining strategies are extensively analyzed in the literature with notable works such as [4, 6, 5]. In this work, we consider a sub-group of strategies from all possible states and action space, which we call $L$-stubborn [^1] In this paper, $L$-stubborn mining is different than the lead-stubbornness (L-stubbornness) defined in [6]. There, “L” stands for the word “lead.” Here, $L$ refers to an integer which is the length of the adversarial chain with respect to a common ancestor block between the honest and the adversarial chains. mining. Here, we explain how $L$-stubborn mining strategy works explicitly and later in Appendix B.3, we present a more formal description of state and action spaces for the same strategy.
To describe the strategy, let us assume that all participants start with the same chain, which we call the offset chain. For the sake of simplicity, we denote the height of the offset chain as zero. Both the adversary and the honest miners mine on top of the offset chain. We denote the highest adversarial block height as $A$ and the chain starting from the offset and ending in the adversarial block at height $A$ as the $A$-chain. We denote the highest honest block height as $H$ and the chain starting from the offset and ending in the honest block at height $H$ as the $H$-chain.
Starting from the offset chain:
@@
    If the adversary is the first to mine a block on top of the offset chain, i.e., at height $1$, it keeps the block private and continues to mine privately on top of its block. At any time later, if the honest miners mine a block at height $h\leq A<L$, the adversary releases its block at the same height $h$ immediately which we call matching. In accordance with the network model defined earlier, $\gamma$ proportion of the honest miners accept and mine on top of the adversarial block instead of the honest block.
@@
An attack cycle is a process that starts when an offset chain is defined, and ends when a new offset chain is defined. Note that, offset chain is redefined whenever an adopt or override action (event) happens, which we call an offset event in general. If during an attack cycle, the length of the $A$-chain reaches $L$, i.e., $A=L$, then we call the attack successful. On the other hand, if the length of the $H$-chain surpasses the length of the $A$-chain before $A$ reaches $L$, we call the attack cycle unsuccessful.
Here, we consider an example that is critical in understanding the model and the subsequent analysis in later parts of the paper. Let us assume $L=7$ and currently $A=3$, $H=0$ at an attack cycle where the private $A$-chain is $A=\mathsf{B^{\prime}}|\mathsf{B_{2}^{\prime}}|\mathsf{B_{3}^{\prime}}$ where ‘$|$’ represents two consecutive blocks chained/concatenated. Assume the next block mined is an honest block $\mathsf{B}$ which is at the same height as $\mathsf{B^{\prime}}$ since honest miners are not aware of any block at that height. As soon as $\mathsf{B}$ is mined, the adversary matches it, i.e., releases $\mathsf{B^{\prime}}$. At this point, $A=3$, $H=1$ and $\gamma$ fraction of honest miners are mining on top of $\mathsf{B^{\prime}}$ whereas the rest of the honest miners mine on top of $\mathsf{B}$. If the next block is honest, i.e., $\mathsf{B_{2}}$, then $A=3$, $H=2$ and the adversary matches $\mathsf{B_{2}}$ with $\mathsf{B_{2}^{\prime}}$ but there are two cases regarding the prefix of $\mathsf{B_{2}}$. Either $H=\mathsf{B}|\mathsf{B_{2}}$ with probability $1-\gamma$ or $H=\mathsf{B^{\prime}}|\mathsf{B_{2}}$ with probability $\gamma$. Hence, during an attack cycle, the prefix of $H$-chain can change and contain some number of adversarial blocks. [^2] A natural question is, if $A=\mathsf{B^{\prime}}|\mathsf{B_{2}^{\prime}}|\mathsf{B_{3}^{\prime}}$ and $H=\mathsf{B^{\prime}}|\mathsf{B_{2}}$, then why not redefine offset chain with $\mathsf{B^{\prime}}$ at its tip since it is accepted by everyone. If we do so, then $A=2$ with blocks $A=\mathsf{B_{2}^{\prime}}|\mathsf{B_{3}^{\prime}}$ and $H=1$ with block $\mathsf{B_{2}}$. However, this renders attack cycles as dependent and we lose track of the double-spending between blocks $\mathsf{B}$ and $\mathsf{B^{\prime}}$. This is an essential design choice for tractability purposes, see Section 7.2 for possible extensions.
3.1.1 An Example: $4$-Stubborn Mining
In Figure 1, we display the sample paths of $4$-stubborn mining, i.e., $L=4$. The $y$-axis and $x$-axis represent $A$ and $H$, i.e., the length of $A$-chain and $H$-chain, respectively. An attack starts from point $(0,0)$ and arrivals of adversarial (honest resp.) blocks are represented with light-green (orange resp.) arrows. A red arrow shows that the arrival of honest block has ended the cycle. A blue (red resp.) cross represents the endpoint of a successful (unsuccessful resp.) attack cycle. A dark-green arrow represents the arrival of an adversarial block that guarantees a successful attack cycle. The dark-green crosses represent the first time cycles are guaranteed to be successful and not ended yet. The cycles that pass through the dark-green crosses eventually end with an honest arrival in blue cross points (x,x1)(x,x-1) where $x>4$, i.e., when the difference of chain lengths reduces to $1$. The notion of $P[n,m]$ at $(A,H)=(n,m)$ displayed in Figure 1 represents the number of distinct paths an attack cycle can take to reach the point $(A,H)=(n,m)$. In Appendix A we give the formula for $P[n,m]$ as well as explain the reasoning behind.
Note that there is a third dimension of Figure 1, which tracks the number of adversarial blocks in the prefix of $H$-chain. Further, the dynamics of this dimension is Markovian since each honest arrival either increases the previous value to H1H-1 (with probability $\gamma$) or stays the same (with probability $1-\gamma$) and does not depend on anything else. This Markovian structure implies that we can separate the third dimension and treat it independently which simplifies our analysis, hence, we do not plot it; see Appendix B.3 for more details.
Figure 1: Sample paths of $4$-stubborn mining.
3.1.2 Relation to the Existing Works
@@
For $L=1$ the strategy reduces to honest mining strategy, whereas for $L=2$ the strategy reduces to the celebrated selfish mining attack of [4]. On the other hand, for $L=\infty$ the strategy becomes the equal fork stubborn mining strategy described in [6] and analyzed in closed-form with Catalan numbers in [37]. In this paper, we are interested in giving a method for calculating the revenue ratios for any given $L$ as well as deriving a method to find the best possible $L$ given $\alpha$ and $\gamma$. The resulting strategy does not allow any trail stubbornness discussed in [6], i.e., the adversary accepts the honest chain whenever it falls behind. Hence, we can only claim the optimality of the strategy under action spaces with independent attack cycles that allow no-trail stubbornness. We argue that the parameter $L$ represents the level of stubbornness in selfish mining and our analysis provides the optimal level of stubbornness as well as the maximum stubbornness such that the attack is still more profitable than honest mining.
@@
In [5], the authors use an undiscounted average reward MDP to provide $\epsilon$-optimal policies. The authors of [5] further argue about combining double-spending attack with selfish mining attack, however, they do not provide any concrete method or analysis. Later, in [8], the same authors provide an MDP that accounts for rewards of double-spending attack. However, the results and strategies provided in [8, 5, 7] are complicated with no easy-to-follow structure. On the other hand, $L$-stubborn mining that we consider in this paper is a policy which results in higher mining rewards than honest mining and we provide an efficient algorithm to obtain the optimal block mining revenue through optimizing $L$, which has strong implications on double-spending strategies with a clear and easy-to-follow analysis under $k$-confirmation rule, which we explain next.
3.2 Double-Spend Events
@@
Here, we rigorously define the events related to a double-spending attack.
@@
($\bm{k}$-Confirmation rule) A block is said to be confirmed according to the $k$-confirmation rule if it is part of a longest public chain and $k$-deep, i.e., there are at least k1k-1 public blocks mined on top of it in this longest chain.
Definition 2
(Double-Spend Events [^3] It is worth mentioning that according to the adversarial action space we consider in this paper, the occurrence of the event Service implies that $\mathsf{B^{\prime}}$ is never confirmed, i.e., only $\mathsf{B}$ is confirmed, although an unrestricted state/action space does not exclude such an instance. Similarly, Move-Funds implies that only $\mathsf{B^{\prime}}$ is confirmed and Double-Spending implies that first $\mathsf{B}$ is confirmed but then $\mathsf{B^{\prime}}$ is confirmed and makes into the offset chain. These three events exhaust all the possibilities regarding the confirmation of block $\mathsf{B}$ and $\mathsf{B^{\prime}}$ according to the attacks we consider in this paper.) Consider an honest block $\mathsf{B}$ and an adversarial block $\mathsf{B^{\prime}}$ at the same height [^4] The adversarial block $\mathsf{B^{\prime}}$ is assumed to contain a transaction where the adversary simply moves some funds between its sybil accounts, whereas the honest block $\mathsf{B^{\prime}}$ is assumed to contain a conflicting transaction where the adversary pays a merchant for some service using the same funds. As observed by [5], Service or Move-Funds events should not be seen as a loss from the perspective of the adversary.. We consider three disjoint events regarding the confirmation of $\mathsf{B}$ and $\mathsf{B^{\prime}}$:
@@
    Service If $\mathsf{B}$ is confirmed and makes it into the offset chain, we call the event Service.
@@
    Move-Funds If $\mathsf{B}$ is never confirmed, we call the event Move-Funds.
@@
    Double-Spending If $\mathsf{B}$ is confirmed but does not make it into the offset chain, we call the event Double-Spending.
Notice the relation between $L$-stubborn mining and $k$-confirmation rule. If $(k+1)$-stubborn mining attack is more profitable than honest mining, at every attack cycle, the adversary gets a shot at double-spending at no-cost while implementing the $(k+1)$-stubborn mining strategy.
Before the offset event happens, if $A>k$, this implies that at some point $H=k$ and the adversary will release $A$-chain later on. Depending on whether $H$-chain contains $\mathsf{B^{\prime}}$ or $\mathsf{B}$, either Move-Funds or Double-Spending happens. On the other hand, before the offset event happens, if $A<H\leq k$, depending on whether $H$-chain contains $\mathsf{B^{\prime}}$ or $\mathsf{B}$, either Move-Funds or Service happens. If an attack cycle ends with Move-Funds event, later, the adversary can try double-spending again using the same funds. One could also rescale the probabilities of Double-Spending and Service by excluding Move-Funds event from the perspective of the actual funds associated with $\mathsf{B}$ and $\mathsf{B^{\prime}}$ since eventually the fund is going to result in Service or Double-Spending. We omit the rescaling in our calculations and numerical results, as rescaling is straightforward to do and we would like to present the exact probability of Double-Spending event in each attack cycle.
We also note that if we consider transactions instead of blocks, then the time $\mathsf{tx}$ is released in the network is another important issue while considering double-spending attacks. In general, it can be shown that a double-spending attack can be successful with probability $1$ if the adversary is allowed to pick the time $\mathsf{tx}$ is released, since it can build an arbitrarily high pre-mining gain [8]. The pre-mining gain refers to the difference between the private chain of the adversary and the public honest chain, in general [8]. The approach we take here is more conservative since we only consider multiple block confirmations at the same height as double-spending, hence the pre-mining gain is 0.
@@ 3.3 $S$-Stealth Mining
Notice the difference between Nakamoto private attack and $(k+1)$-stubborn mining. The first major difference is $(k+1)$-stubborn mining gives up the attack when $A$-chain falls behind $H$-chain, i.e., no-trailing behind is allowed. Another major difference is, during a cycle, $(k+1)$-stubborn mining matches each honest block with an adversarial one which can expose the attack as two chains with conflicting blocks are competing openly in the network, arousing the merchant’s suspicion. The matching also decreases Double-Spending probability while increasing Move-Funds probability, which we show later.
We address some of these issues by modifying $L$-stubborn mining in a way that only allows matching if both double-spending and the risk of losing mining revenue are imminent, i.e., H=A=k=L1H=A=k=L-1. We call this attack $S$-stealth mining, which we investigate in Section 5. Essentially, $S$-stealth mining strategy is the same strategy as $L$-stubborn mining strategy except that the matching action only happens when H=A=S1H=A=S-1. We note that $S$-stealth mining does not allow any trailing behind either, hence, there is still a difference between a private attack and $S$-stealth mining. Note that, similar to $L$ of $L$-stubborn mining strategy, we can view $S$ of $S$-stealth mining strategy as a parameter of stubbornness. We present a more formal description of state and action spaces for $S$-stealth mining in Appendix B.3.